# The Cosmic Dance: Observational Detection of Coherent Spin in Galaxy Clusters

### v1

@@ 3.2 Radial Profile of Cluster Spin
To further quantify the rotational motion suggested by the spatial segregation in the left panel, the middle panel of Figure 4 presents the cluster rotation speed at given radius as a function of normalized distance from the cluster center. Here, each point represents the measured rotation speed of galaxies within a given distance bin. Blue and red points correspond to galaxies on opposite sides of the rotation axis, as indicated by the background shading. The solid lines show the median rotation speed in bins of normalized distance, and the shaded bands indicate the associated uncertainties. Notably, the rotation speed increases with radius. The uncertainties also become larger at greater radius, primarily because the number of galaxies decreases with increasing distance from the cluster center.
In the right panel of Figure 4, we examine the rotation speed $\mu$ of galaxy clusters, defined in Equ. 3, which corresponds to the upper $x$-axis in Figure 3, as a function of the cluster virial mass. Each data point represents an individual cluster, with the color indicating the size of the cluster’s virial radius, ranging from smaller (blue) to larger (red) clusters. The solid line shows the mean distribution, while the dashed lines indicate the $1\sigma$ statistical uncertainty range. It can be seen that as the cluster mass increases, the rotation speed also increases, rising from approximately $\sim 330{\,\rm km\,s^{-1}}$ at $M_{\rm vir}\sim 10^{14}M_{\odot}$ to about $\sim 800{\,\rm km\,s^{-1}}$ at $M_{\rm vir}\sim 10^{15}M_{\odot}$. This positive correlation suggests that more massive clusters tend to exhibit stronger rotational motion.
@@
Fig4a.pngFig4b.pngFig4c.png
Figure 4: Left: Projected distribution of galaxies within clusters, where the $x$- and $y$-axes represent normalized coordinates ($\Delta x/d_{\mathrm{max}}$, $\Delta y/d_{\mathrm{max}}$) centered on the cluster center, with $d_{\mathrm{max}}$ defined as the distance to the most distant member galaxy in each cluster. Each point is color-coded by its redshift deviation $\Delta z$ from the cluster mean; redder (bluer) points indicate galaxies moving away (approaching). The black vertical line marks the rotation axis, and the arrow shows the inferred spin direction. All clusters are rotated to align their spin axes vertically. Middle: Cluster rotation speed as a function of normalized distance to the center. Each point shows the measured rotation speed in a distance bin, with blue and red points for galaxies on opposite sides of the rotation axis. Solid lines and shaded regions indicate the median and uncertainty. Distances in the receding (approaching) region are shown in red (blue) and assigned positive (negative) values. Right: The relation between circular velocity $\log\mu$ (km s-1) and cluster mass, $\log(M_{\rm 200}/M_{\odot})$. Each data point represents a single cluster, color-coded by its physical radius $R_{\rm cluster}$ as indicated by the colorbar. The black solid line shows the median relation, while the grey dashed lines represent the $1\sigma$ scatter in each mass bin.
@@ 3.3 Alignment of Cluster Spin
Figure 5 presents the results of this alignment analysis. In the left panel, we show the distribution of the angle between the cluster spin axis and the spin of the central galaxy. Here, the spin of the central galaxy is calculated following the methods of Lee & Erdogdu (2007); Varela et al. (2012); Kraljic et al. (2021); Zhang et al. (2015); Wang et al. (2025a), using the observed position angle, inclination, and sky coordinates. The blue line represents the full cluster sample, while the red line corresponds to clusters with more than 10 member galaxies ($N_{\mathrm{gal}}>10$). Both samples exhibit a clear tendency for parallel alignment between the cluster spin and the central galaxy spin, with the alignment signal being notably stronger in the richer ($N_{\mathrm{gal}}>10$) subsample.
The right panel displays the distribution of the angle between the cluster spin axis and the orientation of the nearest filament. To determine the filament orientation, we identify the closest filament to each cluster using the filament catalog used in Wang et al. (2021) with Bisous model (Tempel et al., 2014a), and adopt the local filament orientation as the reference axis. The results show a preference for perpendicular signal between the cluster spin and the filament orientation, with this signal being more pronounced in clusters with $N_{\mathrm{gal}}>10$. This finding is consistent with recent observational (Rong et al., 2025) and simulation (Wang et al., 2025) results.
@@ 4 Summary
In this study, we investigated the spin properties of galaxy clusters using the galaxy group catalog and member galaxy data from the Sloan Digital Sky Survey (SDSS) as presented by Tempel et al. (2017). Our cluster sample consists of 2,170 systems with $M>10^{14}\,M_{\odot}$, selected directly from the group catalog. We quantified the spin of each cluster by searching for the orientation in the projected plane that maximizes the redshift difference between two hemispheres of member galaxies, and assessed the statistical significance of the detected spin signals using Monte Carlo randomizations.
Our results provide strong statistical evidence for the presence of coherent rotational motion in galaxy clusters. The observed distribution of the maximum redshift difference, $\Delta Z_{\rm max}$, between member galaxies on opposite sides of the projected rotation axis, shows a significant excess compared to randomized control samples, with a confidence level reaching nearly $150\sigma$ at $\sim 380{\,\rm km\,s^{-1}}$. This signal is even more pronounced in richer clusters with more than 10 member galaxies. Stacked visualizations further confirm that the projected rotation axes effectively separate redshifted and blueshifted galaxy populations within clusters.
@@ 5 Discussion
The origin of cluster angular momentum is intimately connected to the large-scale structure of the Universe (Vitvitska et al., 2002). It is widely recognized that the majority of the mass accreted by galaxy clusters is funneled through the cosmic filaments (Libeskind et al., 2013, 2014; Kang & Wang, 2015) that connect to them, making these filaments the dominant channels for the buildup of cluster angular momentum. In the framework of hierarchical structure formation, galaxy clusters correspond to the most massive dark matter halos, whose angular momentum is thought to arise primarily from the anisotropic infall of matter along these filaments. Numerical simulations have consistently shown that the angular momentum vectors of massive halos tend to be oriented perpendicular (Wang & Kang, 2017, 2018; Ganeshaiah Veena et al., 2018, 2019, 2021) to the direction of the nearest filament. When using member galaxies as tracers to estimate the angular momentum of the underlying dark matter halo (Wang et al., 2025), the resulting distribution is expected to reflect that of the halo itself.
Our results, together with recent observational findings (Rong et al., 2025), indicate that the spin axes of clusters are preferentially perpendicular to the orientation of the nearest filament. As shown in the right panel of Figure 5 and also reported by Rong et al. (2025) and Wang et al. (2025), this perpendicular signal between cluster spin and nearby filament is present but relatively weak. This is consistent with a scenario in which the infall of matter along filaments imparts a net rotational motion to the cluster, establishing a preferred spin-filament alignment. In addition to filamentary accretion (Wang & Kang, 2018), clusters may also acquire angular momentum through the accretion of more diffuse material or isolated galaxies from the surrounding environment (Welker et al., 2014). The combined effects of these processes shape the observed spin properties of galaxy clusters and underscore the complex interplay between cluster growth and the cosmic web.
@@
We thank Prof. Xi Kang (Zhejiang University) and Prof. Tao Wang (Nanjing University) for insightful discussions and comments. PW acknowledge the financial support from the NSFC (No.12473009), and also sponsored by Shanghai Rising-Star Program (No.24QA2711100). This work is supported by the China Manned Space Program with grant no. CMS-CSST-2025-A03. Y.R. acknowledges supports from the CAS Pioneer Hundred Talents Program (Category B), the NSFC grant 12273037, the USTC Research Funds of the Double First-Class Initiative.
@@ Appendix A Ambiguity in the Determination of the Projected Rotation Axis
The right panel of Figure A1 presents the normalized cumulative distribution function (CDF) of $\theta_{\rm max}$ for the entire cluster sample (blue) and for clusters with richness greater than 10 (orange). For the full sample, approximately 55% of clusters have $\theta_{\rm max}$ less than $20^{\circ}$, indicating that in many cases the rotation axis is relatively well constrained. For the richer subsample with more than 10 member galaxies, this fraction increases to about 85%, demonstrating that the ambiguity in determining $\theta_{\rm max}$ is significantly reduced when more member galaxies are available. This highlights the importance of sample richness in reliably identifying the projected rotation axis and minimizing uncertainties due to sparse or asymmetric galaxy distributions.

### v3

@@ 3.2 Radial Profile of Cluster Spin
To further quantify the rotational motion suggested by the spatial segregation in the left panel, the middle panel of Figure 4 presents the cluster rotation speed at given radius as a function of normalized distance from the cluster center. Here, Each point represents the measured galaxy rotation speed at a given distance from the cluster center, normalized by $d_{\rm max}$. Blue and red points correspond to galaxies on opposite sides of the rotation axis. Notably, the rotation speed increases with radius. The scatter becomes larger at greater radius, primarily because the number of galaxies decreases with increasing distance from the cluster center.
In the right panel of Figure 4, we examine the rotation speed $\mu$ of galaxy clusters, defined in Equ. 3, which corresponds to the upper $x$-axis in Figure 3, as a function of the cluster virial mass. Each data point represents an individual cluster, with the color indicating the number galaxies of the clusters, ranging from smaller (blue) to larger (red) clusters. The solid line shows the mean distribution, while the dashed lines indicate the $1\sigma$ statistical uncertainty range. It can be seen that as the cluster mass increases, the rotation speed rises from $\sim 360\,\mathrm{km\,s^{-1}}$ at $M_{\rm vir}\sim 10^{14}\,M_{\odot}$ to $\sim 693)\,\mathrm{km\,s^{-1}}$ at $M_{\rm vir}\sim 10^{15}\,M_{\odot}$. This positive correlation suggests that more massive clusters tend to exhibit stronger rotational motion. The distribution of the red data points (corresponding to larger values of $N_{\mathrm{gal}}$) generally follows the same trend as that of the blue data points (with smaller $N_{\mathrm{gal}}$), but is more tightly clustered around the black solid line. This indicates that, although the high-$N_{\mathrm{gal}}$ systems are fewer in number, they dominate the relation between rotation speed $\mu$ and cluster mass.
@@
s1_f4a.pngs1_f4b.pngs1_f4c.png
Figure 4: Left: Projected distribution of galaxies within clusters, where the $x$- and $y$-axes represent normalized coordinates ($\Delta x/d_{\mathrm{max}}$, $\Delta y/d_{\mathrm{max}}$) centered on the cluster center, with $d_{\mathrm{max}}$ defined as the distance to the most distant member galaxy in each cluster. Each point is color-coded by its redshift deviation $\Delta z$ from the cluster mean; redder (bluer) points indicate galaxies moving away (approaching). The black vertical line marks the rotation axis, and the arrow shows the inferred spin direction. All clusters are rotated to align their spin axes vertically. Middle: Cluster rotation speed as a function of normalized distance to the center. Each point shows the measured rotation speed of a member galaxy, with blue and red points for galaxies on opposite sides of the rotation axis. Distances in the receding (approaching) region are shown in red (blue) and assigned positive (negative) values. Right: The relation between circular velocity $\log\mu$ (km s-1) and cluster mass, $\log(M_{\rm 200}/M_{\odot})$. Each data point represents a single cluster, color-coded by its richness, $N_{gal}$, as indicated by the colorbar. The black solid line shows the median relation, while the grey dashed lines represent the $1\sigma$ scatter in each mass bin.
@@ 3.3 Alignment of Cluster Spin
Early work on the alignment of central cluster galaxies and their host clusters has been extensively studided. The observational support for central galaxy alignment is solid and uncontested, Most studies determine the presence and dependence of alignment by comparing the major axis of the central galaxy’s shape (elliptical fit) with that of the host cluster (calculated from the distribution of member galaxies using inertial momentum or by fitting the X-ray gas distribution)(e.g., Sastry 1968; Binggeli 1982; Niederste-Ostholt et al. 2010; Huang et al. 2016; West et al. 2017; Wittman et al. 2019; Yuan & Wen 2022). By specifying the rotation axes of the clusters in our method, we can further supplement and validate the alignment between the central galaxy and the host cluster.
Figure 5 presents the results of this alignment analysis. In the left panel, we show the distribution of the angle between the cluster spin axis and the spin of the central galaxy. Here, the spin of the central galaxy is calculated following the methods of Lee & Erdogdu (2007); Varela et al. (2012); Kraljic et al. (2021); Zhang et al. (2015); Wang et al. (2025a), using the observed position angle, inclination, and sky coordinates. The blue line represents the full cluster sample, while the red line corresponds to clusters with more than 10 member galaxies ($N_{\mathrm{gal}}>10$). Both samples exhibit a clear tendency for parallel alignment between the cluster spin and the central galaxy spin, with the alignment signal being notably stronger in the richer ($N_{\mathrm{gal}}>10$) subsample. Recently, Wang (2025) investigated in detail the alignment between the spin axes of galaxy clusters and their central galaxies, finding that the alignment strength depends on both the cluster mass and the mass of the central galaxy. Our results are consistent with their findings.
The right panel displays the distribution of the angle between the cluster spin axis and the orientation of the nearest filament. To determine the filament orientation, we identify the closest filament to each cluster using the filament catalog used in Wang et al. (2021) with Bisous model (Tempel et al., 2014a), and adopt the local filament orientation as the reference axis. The results show a preference for perpendicular signal between the cluster spin and the filament orientation, with this signal being more pronounced in clusters with $N_{\mathrm{gal}}>10$. This finding is consistent with recent observational (Rong et al., 2025) and simulation (Wang et al., 2025c) results.
@@ 4 Summary
In this study, we investigated the spin properties of galaxy clusters using FoF-based group catalog (Tempel et al., 2017) and the halo-based group catalog (Yang et al., 2012). Our cluster samples consists of 2,170 and 1329 systems with $M>10^{14}\,M_{\odot}$, respectively. We quantified the spin of each cluster by searching for the orientation in the projected plane that maximizes the redshift difference between two hemispheres of member galaxies, and assessed the statistical significance of the detected spin signals using Monte Carlo randomizations.
Our results provide strong statistical evidence for the presence of coherent rotational motion in galaxy clusters. The observed distribution of the maximum redshift difference, $\Delta Z_{\rm max}$, between member galaxies on opposite sides of the projected rotation axis, shows a significant excess compared to randomized control samples, with a confidence level reaching its peak at $\sim 380{\,\rm km\,s^{-1}}$ for Sample-1 and at $\sim 300{\,\rm km\,s^{-1}}$ for Sample-2. This signal is even more pronounced in richer clusters with more than 10 member galaxies. Stacked visualizations further confirm that the projected rotation axes effectively separate redshifted and blueshifted galaxy populations within clusters.
@@ 5 Discussion
The origin of cluster angular momentum is intimately connected to the large-scale structure of the Universe (Vitvitska et al., 2002). It is widely recognized that the majority of the mass accreted by galaxy clusters is funneled through the cosmic filaments (Libeskind et al., 2013, 2014; Kang & Wang, 2015) that connect to them, making these filaments the dominant channels for the buildup of cluster angular momentum. In the framework of hierarchical structure formation, galaxy clusters correspond to the most massive dark matter halos, whose angular momentum is thought to arise primarily from the anisotropic infall of matter along these filaments. Numerical simulations have consistently shown that the angular momentum vectors of massive halos tend to be oriented perpendicular (Wang & Kang, 2017, 2018; Ganeshaiah Veena et al., 2018, 2019, 2021) to the direction of the nearest filament. When using member galaxies as tracers to estimate the angular momentum of the underlying dark matter halo (Wang et al., 2025c), the resulting distribution is expected to reflect that of the halo itself.
Our results, together with recent observational findings (Rong et al., 2025), indicate that the spin axes of clusters are preferentially perpendicular to the orientation of the nearest filament. As shown in the right panel of Figure 5 and also reported by Rong et al. (2025) and Wang et al. (2025c), this perpendicular signal between cluster spin and nearby filament is present but relatively weak. This is consistent with a scenario in which the infall of matter along filaments imparts a net rotational motion to the cluster, establishing a preferred spin-filament alignment. In addition to filamentary accretion (Wang & Kang, 2018), clusters may also acquire angular momentum through the accretion of more diffuse material or isolated galaxies from the surrounding environment (Welker et al., 2014). The combined effects of these processes shape the observed spin properties of galaxy clusters and underscore the complex interplay between cluster growth and the cosmic web.
@@
We thank Prof. Xi Kang (Zhejiang University) and Prof. Tao Wang (Nanjing University) for insightful discussions and comments. PW acknowledge the financial support from the NSFC (No.12473009), and also sponsored by Shanghai Rising-Star Program (No.24QA2711100). This work is supported by the China Manned Space Program with grant no. CMS-CSST-2025-A03. Y.R. acknowledges supports from the CAS Pioneer Hundred Talents Program (Category B), the NSFC grants 12522302 and 12273037, and the USTC Research Funds of the Double First-Class Initiative. This work is also supported by the China Manned Space Program with grant no. CMS-CSST-2025-A06 and CMS-CSST-2025-A08. M.B. acknowledges support by the National Natural Science Foundation of China, NSFC grant No. 12303009.
@@ Appendix A Ambiguity in the Determination of the Projected Rotation Axis
The right panel of Figure A1 presents the normalized cumulative distribution function (CDF) of $\Delta\theta_{\rm max}$ for the entire cluster sample (blue) and for clusters with richness greater than 10 (orange). For the full sample, approximately 55% of clusters have $\Delta\theta_{\rm max}$ less than $20^{\circ}$, indicating that in many cases the rotation axis is relatively well constrained. For the richer subsample with more than 10 member galaxies, this fraction increases to about 85%, demonstrating that the ambiguity in determining $\Delta\theta_{\rm max}$ is significantly reduced when more member galaxies are available. This highlights the importance of sample richness in reliably identifying the projected rotation axis and minimizing uncertainties due to sparse or asymmetric galaxy distributions.
@@
Appendix B Distributions of N and $\Delta Z_{\rm max}$, and their relationship
This appendix provides additional summaries of cluster richness and the redshift contrast metric to complement the main text. Figure A2 presents the distribution of the number of member galaxies per cluster ($N_{gal}$, left panel), the distribution of $\Delta Z_{\max}$ ( middle panel), and $\Delta Z_{\max}$ as a function of $N_{gal}$ (right panel) for both cluster samples. The sample-1 is dominated by sparse systems, with the majority of clusters containing fewer than 20 member galaxies,while Sample-2 includes clusters with a larger number of member galaxies. The distribution of $\Delta Z_{\max}$ is unimodal and peaks near 0.003 for Sample-1, whereas for Sample-2 it is smaller, peaking around 0.002. Consistent with the visual impression in the right panel, $\Delta Z_{\max}$ does not exhibit a strong dependence on cluster richness ($N_{\mathrm{gal}}$), although richer clusters show slightly larger scatter, and for Sample-2 there is a weak increase of $\Delta Z_{\max}$ with $N_{\mathrm{gal}}$.
[…]
s12_a1.pngs12_a2.pngs12_a3.png
Figure A2: Distributions of the basic properties of the galaxy cluster samples used for both Sample-1 and Sample-2. Left: distribution of the number of galaxies per cluster. Middle: distribution of $\Delta Z_{\max}$. Right: $\Delta Z_{\max}$ versus the number of galaxies per cluster.